### v1

@@ 2.1 $\delta$-locality
    ($\ell_{1}$-condition) $H(x)=-\nabla^{2}\log\pi(x)$ is $c$-unifomly diagonal block dominant, i.e. there exists a matrix $M\in\mathbb{R}_{+}^{b\times b}$ s.t. $\forall i,j\in[b],i\neq j$,
@@
    where $H_{ij}(x)$ denotes the $(i,j)$-th subblock of $H(x)$. Here $\delta$ can be taken as $\delta=c^{-1}$.
  2. (2)
    ($\ell_{2}$-condition) $H(x)=-\nabla^{2}\log\pi(x)$ has uniform block bandwidth $B>0$, and uniform lower and upper bounded $m$ and $M$ where $0<m\leq M<\infty$, i.e.
    $\forall|i-j|>B,\quad H_{ij}(x)=0;\quad mI\preceq H(x)\preceq MI,$
    Here $\delta$ can be taken as $\delta=\frac{2BM}{m^{2}}$.
Proofs can be found in Section 6.1.
Remarks 2.3.
(1) Note that the two conditions above both require that $\pi$ is log-concave. To see that the $\ell_{1}$-condition implies log-concavity, denote $M_{ij}^{\prime}:=2M_{ii}\mathbf{1}_{i=j}-M_{ij}$, then $M^{\prime}$ is $c$-diagonal dominant. Gerŝgorin discs theorem [12] then implies the smallest eigenvalue of $M^{\prime}$ is lower bounded by $c$, and thus
$u^{\rm T}H(x)u=~{}$ $\sum_{i,j}u_{i}^{\rm T}H_{ij}(x)u_{j}\geq\sum_{i}M_{ii}\left|u_{i}\right|^{2}-\sum_{i\neq j}M_{ij}\left|u_{i}\right|\left|u_{j}\right|$
@@
(2) The two conditions characterize the locality of the distribution in different ways. The $\ell_{1}$-condition uses diagonal dominance of the Hessian, which implies that any block is mostly influenced by itself. The $\ell_{2}$-condition directly uses bandedness of Hessian to characterize the localized structure.
2.2 $\ell_{\infty}$-approximation error
@@
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ is $\delta$-localized, then the marginal $W_{1}$-distance of $\pi,\pi^{\prime}$ satisfies
$\max_{i\in[b]}W_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)},$ (2.3)
where $\pi_{i}$ and $\pi_{i}^{\prime}$ are the marginals of $\pi$ and $\pi^{\prime}$ on $x_{i}$ respectively.
@@
$W_{1}(\pi_{i},\pi_{i}^{\prime})=\sup_{\phi_{0}\in{\rm Lip}_{1}}\int\phi_{0}(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i},$
where ${\rm Lip}_{1}$ denotes the $1$-Lipchitz function class. Fix any $\phi_{0}\in{\rm Lip}_{1}$. Denote
$\phi(x_{i})=\phi_{0}(x_{i})-\int\phi_{0}(x_{i})\pi^{\prime}(x_{i})\mathrm{d}x_{i}.$
Then
$\int\phi_{0}(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}=~{}$ $\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}$
@@
Here we use $\pi(x_{i})=\int\pi(x_{i},x_{-i})\mathrm{d}x_{-i}$. Let $u(x)$ solve the Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{i}).$
Now by integration by parts,
$\int\phi(x_{i})\pi(x)\mathrm{d}x=~{}$ $\int\left(\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)\right)\pi(x)\mathrm{d}x$
@@
$\int\phi(x_{i})\pi(x)\mathrm{d}x\leq~{}$ $\sum_{j=1}^{b}\left\|\nabla_{j}u\right\|_{L^{\infty}}\int\left|\nabla_{j}\log\pi^{\prime}(x)-\nabla_{j}\log\pi(x)\right|\pi(x)\mathrm{d}x$
@@
$\max_{i}W_{1}(\pi_{i},\pi_{i}^{\prime})\leq~{}$ $\max_{i}\max_{\phi_{0}\in{\rm Lip}_{1}}\int\phi_{0}(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}$
@@
2.3 Approximation by sparse graphical models
In this section, we consider a special case of approximating a distribution $\pi$ by an undirected graphical model [16, 27]. The undirected graphical model, also called Markov random field, models the local dependence of the distribution by an undirected graph. It requires that any nonadjacent two components $X_{i},X_{j}$ are conditionally independent given the rest of the components.
Consider an undirected graph $G=(V,E)$ where $|V|=b$. We attach each vertex in $V$ with a block index $i$ uniquely. Denote $N_{j}=\{k\in V:\overline{jk}\in E\}$ as the neighboring vertices of $j$. Here we require $E$ includes all the self-loop in $G$, i.e. $\forall j\in G,\overline{jj}\in E$.
We say a distribution $\pi(x)\in\mathcal{P}(\mathbb{R}^{d})$ has dependence graph $G$ if
$\forall j,k\in V,~{}k\notin N_{j}~{}\Rightarrow~{}x_{j}\perp\!\!\!\perp x_{k}~{}|~{}x_{N_{j}^{-}},$ (2.4)
[…]
$k\notin N_{j}~{}\Rightarrow~{}\nabla_{jk}^{2}\log\pi(x)=0.$
Note it generalizes the banded condition on the Hessian in 2.2 by using the sparse graphical structure. Parallelly, one can show that under certain sparsity condition on $G$, a distribution $\pi$ with dependence graph $G$ is $\delta$-localized.
To specify the sparsity condition, we denote the $q$-th neighbor $N_{j}^{q}~{}(q\in\mathbb{Z}_{+})$ as
$N_{j}^{q}:=\{k\in V:d_{G}(j,k)\leq q\},$
where $d_{G}(j,k)$ is the shortest path distance between $j$ and $k$ in $G$, i.e. the minimum number of edges that must be traversed to go from $j$ to $k$. For convenience, we take $d_{G}(j,j)=0$, and $d_{G}(j,k)=\infty$ if $j,k$ are not connected.
Definition 2.5.
A $s$-sparse graph $G$ with ambient dimension $r\in\mathbb{Z}_{+}$ is such that
$\forall j\in V,\quad|N_{j}^{q}|\leq 1+sq^{r}.$
We also call $G$ a $s$-sparse graph without specifying the ambient dimension.
@@ A $s$-sparse graph $G$ with ambient dimension $r\in\mathbb{Z}_{+}$ is such that
The definition of sparse graph is different from that in graph theory. Roughly speaking, the sparsity $s$ denotes the size of the immediate neighbor, and the dimension $r$ controls the growth rate of the neighbor size with respect to the distance. Note we require it grows at most polynomially to ensure the locality. A typical example of such sparse graph is the lattice model $\mathbb{Z}^{r}$. A naive bound for $\mathbb{Z}^{r}$ is $|N_{j}^{q}|\leq(2q+1)^{r}<1+(3q)^{r}$. So that $\mathbb{Z}^{r}$ is $3^{r}$-sparse with ambient dimension $r$.
Proposition 2.7.
Let $G=(V,E)$ be a $s$-sparse graph with ambient dimension $r$. Suppose $\pi\in\mathcal{P}(\mathbb{R}^{d})$ is a distribution with dependence graph $G$, and satisfies
Then $\pi$ is $\delta$-localized with $\delta=\frac{sr!\kappa^{r}}{m}$ where $\kappa=\frac{M}{m}$. As a consequence, for any $\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$,
$\max_{j\in V}W_{1}(\pi_{j},\pi_{j}^{\prime})\leq\frac{sr!\kappa^{r}}{m}\max_{k\in V}\left\|\nabla_{k}\log\pi-\nabla_{k}\log\pi^{\prime}\right\|_{L^{1}(\pi^{\prime})},$
Proofs can be found in Section 6.2. Note this recovers the second result in 2.2, where the bandedness condition corresponds to a $2B$-sparse graph with ambient dimension $r=1$.
@@ 2.4 Marginal error of multiple blocks
In 2.4, we show that the marginal distance of $\pi^{\prime}$ and a $\delta$-localized distribution $\pi$ on one block $x_{i}$ is dimension independent. Here we extend this result to the case of multiple blocks. This result provides further control on the correlation between different blocks in $\pi^{\prime}$, which may not be captured by 2.4.
@@
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ satisfies any one of the conditions in 2.2 or 2.7, then for any index set $I\subset[b]$, the marginal $W_{1}$-distance of $\pi,\pi^{\prime}$ on $x_{I}:=(x_{i})_{i\in I}$ satisfies
$W_{1}(\pi_{I},\pi_{I}^{\prime})\leq\delta|I|\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.$ (2.5)
Here $\delta$ can be taken as the same constant in 2.2 or 2.7.
Proofs can be found in Section 6.3. The implications of the theorem is two-fold. First, for blocks $i,j$ at large distance, the correlation of $x_{i},x_{j}$ under $\pi$ is neglectable, and $\pi_{i,j}$ are almost independent: $\pi_{i,j}\approx\pi_{i}\pi_{j}$. The above theorem says that $\pi^{\prime}_{i,j}$ is close to $\pi_{i,j}$, so that $x_{i},x_{j}$ are also almost independent under $\pi^{\prime}$. Conversely, for neighboring blocks $i,j$, $\pi^{\prime}_{i,j}$ can capture the correlation between $x_{i},x_{j}$ in $\pi$.
@@ 3 Application I: localized LIS method
where $\pi_{0}(x)$ is the prior distribution, and $l(x)$ is the likelihood function. Sampling from such distributions is a central problem in Bayesian inference. The key in the LIS method is the low rank structure of the likelihood function $l(x)$, making it possible to be approximated by a rigde function
@@
where $P_{r}$ is a projection to a low-dimensional subspace $\mathcal{X}_{r}\subset\mathcal{X}$. The LIS method propose to approximate the target distribution by
$\pi(x)\approx\pi_{r}(x)\propto l_{r}(x_{r})\pi_{0}(x),$
which, with properly chosen informed subspace $\mathcal{X}_{r}$ and ridge approximation of $l(x)$, provides certified approximation error bounds [5]. With such low-rank approximation, the sampling complexity is greatly reduced, and the method has been successfully applied in many high-dimensional problems [4, 3].
@@ 3.1 Localized LIS method
In LIS, the informed subspace is constructed globally, which may not capture the spatial structure in the target distribution. For distributions with sparse dependence graph, a locally defined informed subspace can be a better choice. Based on this observation, we propose to use localized likelihood-informed subspace (LLIS), which incorporates domain decomposition in LIS method. In essence, we decompose the parameters into small blocks, and apply LIS in each block to construct the LLIS.
@@
LLIS construction
We propose to apply the LLIS approximation under a transformed representation so that the prior is a standard Gaussian, i.e. consider the change of variables
@@
Then $\tilde{x}\sim\tilde{\pi}_{0}=\mathcal{N}(0,I_{d})$ if $x\sim\pi_{0}$. Under the transformed representation $\tilde{x}$, we aim to find the LLIS of the product form
@@
where $\widetilde{\mathcal{X}}_{r,j}\subset\mathbb{R}^{d_{j}}$ denotes the subspace of the reduced paramater in block $j$. To determine $\widetilde{\mathcal{X}}_{r,j}$, consider the following diagnostic matrix
@@
Notice compared to the existing LIS methods [5] where only first order information is used, here we also incorporates the second order information in the diagnostic matrix.
@@
where $u_{1},\dots,u_{r_{j}}$ are the eigenvectors of ~$\widetilde{\mathcal{M}}_{jj}$ with the largest $r_{j}$ eigenvalues. Here $r_{j}=\text{\rm rank}(\widetilde{\mathcal{X}}_{r,j})$ is the number of informed directions we choose. It can be determined, for instance, by (see [5])
$r_{j}=\min\left\{1\leq r\leq d_{j}:\sum_{k=1}^{r}\lambda_{k}(\widetilde{\mathcal{M}}_{jj})\geq(1-\epsilon)\text{\rm tr}\left[\widetilde{\mathcal{M}}_{jj}\right]\right\},$
where $\lambda_{k}(\widetilde{\mathcal{M}}_{jj})$ denotes the $k$-th largest eigenvalue of ~$\widetilde{\mathcal{M}}_{jj}$, and $\epsilon\in(0,1)$ is some prescribed threshold.
Given $\widetilde{\mathcal{X}}_{r}$, denote accordingly the complementary space
@@
Here $\tilde{P}_{r,j}$ is the orthogonal projector onto $\widetilde{\mathcal{X}}_{r,j}$, and $\tilde{P}_{\bot,k}$ is its orthogonal complement. Note $x_{r}$ and $x_{\bot}$ are not orthogonal in the usual sense, but it holds that
@@
That is, $x_{r}$ and $x_{\bot}$ are independent under the prior $\pi_{0}=\mathcal{N}(0,C)$.
LLIS approximation
As in the LIS method, we construct the approximate likelihood $l_{r}(x_{r})$ by averaging $l(x)$ in the logarithm regime over the uninformed subspace $\mathcal{X}_{\bot}$, and approximate the target distribution by

### v2

@@ 2.2 Diagonal Dominant Distributions
Consider $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Suppose $H(x):=-\nabla^{2}\log\pi(x)$ is $c$-unifomly diagonal block dominant, i.e. there exists a matrix $M\in\mathbb{R}_{+}^{b\times b}$ s.t. $\forall i,j\in[b],i\neq j$,
@@
where $H_{ij}(x)$ denotes the $(i,j)$-th subblock of $H(x)$. Then $\pi$ is $\delta$-localized with $\delta=c^{-1}$.
Proof can be found in Section 6.3.
Remark 2.2.
The conditions above imply that $\pi$ is log-concave. Denote $M_{ij}^{\prime}:=2M_{ii}\mathbf{1}_{i=j}-M_{ij}$, then $M^{\prime}$ is $c$-diagonal dominant. Gerŝgorin discs theorem [24] implies that the smallest eigenvalue of $M^{\prime}$ is lower bounded by $c$, and thus
$u^{\rm T}H(x)u=~{}$ $\sum_{i,j}u_{i}^{\rm T}H_{ij}(x)u_{j}\geq\sum_{i}M_{ii}\left\|u_{i}\right\|^{2}-\sum_{i\neq j}M_{ij}\left\|u_{i}\right\|\left\|u_{j}\right\|$
@@
Therefore, $\pi$ is $c$-strongly log-concave.
3 Marginal Otto-Villani Inequality
3.1 Marginal Error of One Block
@@
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ is $\delta$-localized, then the marginal $\mathsf{W}_{1}$ distance of $\pi,\pi^{\prime}$ satisfies
$\max_{i\in[b]}\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)},$ (3.1)
where $\pi_{i}$ and $\pi_{i}^{\prime}$ denote the marginals of $\pi$ and $\pi^{\prime}$ on $x_{i}$ respectively.
@@
$\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})=\sup_{\phi\in{\rm Lip}_{1}}\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i},$
where ${\rm Lip}_{1}$ denotes the $1$-Lipchitz function class. Fix any $\phi\in{\rm Lip}_{1}$. Note
$\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}=~{}$ $\int\left(\phi(x_{i})-\mathbb{E}_{\pi_{i}^{\prime}}[\phi(x_{i})]\right)\pi_{i}(x_{i})\mathrm{d}x_{i}$
@@
Here $\mathbb{E}_{\pi_{i}^{\prime}}[\phi(x_{i})]=\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]$ is a constant. Let $u(x)$ solve the marginal Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{i})-\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})].$
By integration by parts,
$\int\left(\phi(x_{i})-\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]\right)\pi(x)\mathrm{d}x=~{}$ $\int\left(\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)\right)\pi(x)\mathrm{d}x$
@@
$\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}\leq~{}$ $\sum_{j=1}^{b}\left\|\nabla_{j}u\right\|_{L^{\infty}}\int\left\|\nabla_{j}\log\pi^{\prime}(x)-\nabla_{j}\log\pi(x)\right\|\pi(x)\mathrm{d}x$
@@
$\max_{i}\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})\leq~{}$ $\max_{i}\max_{\phi\in{\rm Lip}_{1}}\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}$
@@
To control the marginal error, one can directly apply Otto-Villani inequality [34] on the marginals and obtain
$\mathsf{W}_{2}(\pi_{i},\pi_{i}^{\prime})\leq C_{\pi_{i}}\sqrt{\mathbb{E}_{\pi_{i}}\left\|\nabla\log\pi_{i}-\nabla\log\pi_{i}^{\prime}\right\|^{2}}.$
The main issue of this approach is that we often have access only to $\pi$ and $\pi^{\prime}$ but not to their marginals, and evaluating the marginals is often computationally very challenging as it involves integrating out the other components. This makes the inequality less useful in practice.
@@ 3.2 Marginal Error of Multiple Blocks
In 3.1, we show that the marginal distance of $\pi^{\prime}$ and a $\delta$-localized distribution $\pi$ on one block $x_{i}$ is dimension independent. Here we extend this result to the case of multiple blocks.
@@
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ satisfies any one of the conditions in 2.1 or 2.2, then for any index set $I\subset[b]$, the $\mathsf{W}_{1}$ distance of $\pi,\pi^{\prime}$ on the marginal $x_{I}:=(x_{i})_{i\in I}$ satisfies
$\mathsf{W}_{1}(\pi_{I},\pi_{I}^{\prime})\leq\delta|I|\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.$ (3.2)
Here $\delta$ can be taken as the same $\delta$ in 2.1 or 2.2.
Proofs can be found in Section 6.4. 3.2 provides further control on the correlation between different blocks in $\pi^{\prime}$, which cannot be directly derived from 3.1. Roughly speaking, when $\pi$ and $\pi^{\prime}$ are both Gaussians, 3.1 only guarantees that the diagonal blocks of the covariance matrix of $\pi^{\prime}$ are close to those of $\pi$, while 3.2 further guarantees that the off-diagonal blocks are also close.
@@ 4.1 Review of LIS Method
where $\pi_{0}(x)$ is the prior distribution, $l(x)$ is the likelihood function, and $Z$ is the normalizing constant. Sampling from such distributions is a central problem in Bayesian inference.
The key idea in LIS is that the likelihood function $l(x)$ often exhibits a low-rank structure, making it possible to be approximated by a rigde function
@@
where $P_{r}$ is a projection to a low-dimensional likelihood-informed subspace $\mathcal{X}_{r}\subset\mathcal{X}$. This leads to an approximation of the target distribution as
$\pi(x)\approx\pi_{r}(x)\propto l_{r}(x_{r})\pi_{0}(x).$
With a well-chosen subspace $\mathcal{X}_{r}$ and ridge approximation $l_{r}(x)$, LIS provides a certified approximation [12], significantly reducing the sampling complexity. The method has been successfully applied to many high-dimensional Bayesian inference problems [11, 10].
@@ 4.2 Localized LIS Method
In LIS, the informed subspace is constructed globally, which may not capture the spatial structure in the target distribution. For localized distributions, a locally defined informed subspace can be a better choice. Based on this observation, we propose to use localized likelihood-informed subspace (LLIS), which incorporates domain decomposition in LIS method. In essence, we decompose the parameters into small blocks, and apply LIS in each block to construct the LLIS.
@@
4.2.1 Construction of Localized LIS
Suppose $\pi_{0}=\mathcal{N}(0,C)$, and $\pi_{0},\pi$ both have the dependency graph $G=(V,E)$. We will apply the LLIS approximation under the whitened transformation, i.e. consider the change of variables
@@ 4.2.1 Construction of Localized LIS
Note $\tilde{x}\sim\tilde{\pi}_{0}=\mathcal{N}(0,I_{d})$ when $x\sim\pi_{0}$. Under the whitened coordinate $\tilde{x}$, we aim to construct the LLIS of the product form
@@
where $\widetilde{\mathcal{X}}_{r,j}\subset\mathbb{R}^{d_{j}}$ denotes the subspace of the reduced paramater $\tilde{x}_{r,j}$. $\widetilde{\mathcal{X}}_{r,j}$ is determined via the following diagnostic matrix
@@
Notice compared to the existing LIS methods [12], where only the first order information is used, here we also incorporates the second order information in the diagnostic matrix.
@@
where $u_{1},\dots,u_{r_{j}}$ are the top $r_{j}$ eigenvectors of ~$\widetilde{\mathcal{M}}_{jj}$, and $r_{j}=\text{\rm rank}(\widetilde{\mathcal{X}}_{r,j})$ is the number of selected informed directions, which can be determined, for instance, by (see [12])
$r_{j}=\min\left\{1\leq r\leq d_{j}:\sum_{k=1}^{\nu}\lambda_{k}(\widetilde{\mathcal{M}}_{jj})\geq(1-\epsilon)\text{\rm tr}[\widetilde{\mathcal{M}}_{jj}]\right\},$
where $\lambda_{k}(\widetilde{\mathcal{M}}_{jj})$ denotes the $k$-th largest eigenvalue of ~$\widetilde{\mathcal{M}}_{jj}$, and $\epsilon\in(0,1)$ is some prescribed threshold. Given $\widetilde{\mathcal{X}}_{r}$, denote the according complementary space
@@
Here $\tilde{P}_{r,j}$ is the orthogonal projector onto $\widetilde{\mathcal{X}}_{r,j}$, and $\tilde{P}_{\bot,k}$ is its orthogonal complement. Note that $x_{r}$ and $x_{\bot}$ are not orthogonal in the standard $\ell_{2}$ sense, but are orthogonal in the $C$-weighted inner product, i.e.,
@@
This implies that $x_{r}$ and $x_{\bot}$ are independent under the prior $\pi_{0}=\mathcal{N}(0,C)$.
4.2.2 Localized LIS Approximation
As in [12], we construct the approximate likelihood $l_{r}(x_{r})$ by averaging $l(x)$ in the logarithm regime over the uninformed subspace $\mathcal{X}_{\bot}$, and approximate the target distribution by